# SPRINT Enables Interpretable and Ultra-Fast Virtual Screening against Thousands of Proteomes

### v1

@@ Results
To benchmark the usefulness of DTI co-embeddings for marginal property prediction, e.g., predicting the properties of a compound only from its molecular graph, we computed SPRINT DTI co-embeddings for several drug-like compounds [31] and natural products [32]. Concatenating the SPRINT-xs molecule co-embedding to a Morgan fingerprint consistently outperformed an equivalently sized neural network using only the Morgan fingerprint as input (Table S1). However, we observe that using only the SPRINT embedding in these tasks degrades performance, suggesting that SPRINT embeddings can synergistically enhance traditional fingerprints.
@@
Discussion
@@
Vector-based screens are extremely fast, enabling DTI prediction in regimes that would be impossible with structure-based approaches. We propose SPRINT, which improves on prior work using multi-head attention pooling that scales favorably as we increase the number of DTI training tokens. We also show that structure-aware PLMs like SaProt can confer huge performance gains in virtual screening. Interestingly, we find that the SPRINT models that perform best on the LIT-PCBA virtual screening benchmark, with increased negative sampling, have the least interpretable attention maps. We hypothesize that equal weighting of positive and negative drug-target pairs helps the model learn about residues that interact while increasing the amount of negatives dilutes the information of the positive examples. We demonstrate that SPRINT can perform virtual screening at pan-species proteome scales, e.g., for antimicrobials (Fig. S1). Lastly, we find that predicting DTIs via co-embedding is an effective pre-training strategy that enhances simple molecular property prediction (Table S1).
@@
We envision SPRINT as a useful benchmarking tool for protein and molecule encoders. Future work will evaluate other structure-aware PLMs, such as MULAN or S-PLM [33, 34], and pre-trained molecule encoders, like UniMol [35, 36], in the SPRINT framework for DTI prediction.
@@ Appendix A SPRINT recovers known mechanisms of action
Our pan-species protein dataset is comprised of all predicted protein sequences in reference genomes from NCBI within the taxons bacteria, fungus, and human. Each taxon contained 3,379,854, 775,477, and 136,194 protein sequences respectively. We gathered a list of 3,112 natural products [32], as they are known to have a high prior likelihood for antimicrobial activity [37] and are out-of-distribution relative to our MERGED training dataset. We then co-visualized several antimicrobials and drugs with their known protein targets across microbial proteomes (Fig. S1), recovering several known mechanisms of action. The dataset is available to query at https://bit.ly/colab-screen.
Figure S1: UMAP visualization of the binding co-embedding space of drug-like small molecules and their protein targets across bacterial, fungal, and human proteomes. We see that antimicrobial compounds co-localize with regions of the shared latent space that contain human, bacterial, and fungal proteomes.
@@ Appendix B SPRINT co-embeddings improve property prediction
Current antimicrobial and toxicity screening approaches are often formulated as molecular property predictors, framing antimicrobial activity and toxicity to humans as inherent properties of drug molecules [38, 39, 40, 41, 31, 42]. Our results demonstrate that augmenting Morgan fingerprints with SPRINT-xs ligand embeddings consistently outperformed an equivalently sized neural network using Morgan fingerprints alone, when evaluated on both an antibacterial activity dataset [32] and a toxicity dataset [31] (Table S1). We hypothesize that vectorizing the DTI space allows property prediction methods to leverage information about target neighborhoods around a drug, enhancing performance and offering mechanistic explanations for these properties based on likely binding partners. The embeddings from the deeper SPRINT-sm model consistently performed worse than those from the shallow SPRINT-xs model, suggesting that shallow transformations of the Morgan fingerprint work best in this setting. The standalone SPRINT embeddings achieved substantially lower performance than their concatenated counterparts, indicating that SPRINT embeddings may capture complementary molecular features to traditional fingerprints.
Table S1: F1 scores for MLP classification models applied to molecule embedding strategies (mean $\pm$ std). Models trained on Morgan fingerprints used a larger hidden size to match the size of models trained on fingerprints concatenated with embeddings.
@@
[…]
@@ Appendix C Training details
We found that removing the contrastive training on the DUD-E dataset from our reproduced ConPLex model improved its DTI predictive performance. Prior work [44, 45] has shown that the DUD-E dataset [46] has hidden biases that encourage deep learning models to learn drug-only features rather than protein-ligand interactions for determining if a drug is active. Unlike ConPLex, which used DUD-E for contrastive training while using different datasets for binary training, we use the same dataset for binary training and simultaneous contrastive training. We utilize the InfoNCE loss [47, 48], rather than the triplet margin-distance loss employed by ConPLex, to leverage a larger number of negatives per positive example to further constrain the embeddings of the model.
We found that adding InfoNCE to our model decreased performance on all DTI datasets (Table S3), except for the smallest dataset. The InfoNCE loss seems to reduce the overfitting of the SPRINT-sm model on the smaller datasets, but further investigation is needed, as the InfoNCE models performed poorly on downstream tasks like virtual screening. Therefore, we only train our models with the binary cross entropy loss (2) after computing the probability of binding via the sigmoid of the cosine similarity between the protein and drug embeddings (3). Specifically, the loss $\mathcal{L}$ is written as
@@
where the protein, $T^{i}$, and drug, $D^{i}$, have been mapped to the SPRINT co-embedding space as $Z_{t}^{i}$ and $Z_{d}^{i}$, respectively. $Y_{i}\in\{0,1\}$ is a ground-truth label with value 1 if $T^{i}$ and $D^{i}$ are binders or 0 if they are non-binders. The pre-sigmoid scalar value, $\alpha$, is used to expand the range of cosine-similarity to the domain of the sigmoid. We set $\alpha$ to 5.
ConPLex models are trained with the hyperparameters used in the original paper [2]. Our Attention Pooling models are trained with a learning rate of $1\times 10^{-5}$ and a dropout value of 0.05 for 250 epochs, keeping all other hyperparameters the same as ConPLex training. The model checkpoint with the highest validation AUPR during training is evaluated on the test set (Table 1).
To enable efficient training on the MERGED dataset, we featurize the unique proteins and molecules before training, storing their representations in memory-mapped files for quick retrieval using the Lightning Memory-Mapped Database Manager (LMDB) library. We use PyTorch Distributed Data Parallel (DDP) for huge improvements in training speed [49]. To address data imbalance in the binding data, for each epoch, we train using all of the drug-target binding pairs, and subsample an equivalent number of non-binding pairs without replacement. We observed that models trained with more negatives than positives (at a 3:1 ratio), achieved better virtual screening performance, but had less interpretable attention patterns (Tables 1, S2). Models are trained for 20 epochs. All other hyperparameters were kept the same.
Table S2: LIT-PCBA evaluation ablation of negative sampling. ‘1:1’ indicates equal sampling of positive and negative examples during training and ‘3:1’ indicates the preferred model training with 3 negatives sampled for every positive example.
@@
Code and data to reproduce our DTI models are on our GitHub repository https://github.com/abhinadduri/panspecies-dti.
Table S3: AUPR values for DTI prediction with co-embedding models across benchmarks (mean $\pm$ std) ablating the use of InfoNCE. These preliminary results use a pre-sigmoid scalar value of 1 rather than 5.
Model Backbone Pooling BIOSNAP DAVIS BindingDB MERGED
SPRINT-xs (10M) SaProt Attn 0.910 $\pm$ 0.003 0.453 $\pm$ 0.007 0.677 $\pm$ 0.005 0.841 $\pm$ 0.003
SPRINT-xs +InfoNCE SaProt Attn 0.894 $\pm$ 0.002 0.499 $\pm$ 0.006 0.635 $\pm$ 0.010 0.816 $\pm$ 0.008
SPRINT-sm (16M) SaProt Attn 0.850 $\pm$ 0.002 0.425 $\pm$ 0.001 0.560 $\pm$ 0.008 0.844 $\pm$ 0.013
SPRINT-sm +InfoNCE SaProt Attn 0.900 $\pm$ 0.003 0.489 $\pm$ 0.004 0.605 $\pm$ 0.004 0.856 $\pm$ 0.007
Appendix D Investigating the learned aggregation layer
@@
We first analyzed the attention patterns of each head in the learned aggregation layer to determine if any of the heads were selective for binding residues. We calculate the attention scores for each residue in the protein and compute the mean attention value for the binding residues and the non-binding residues. Figure 2 shows the average weight of the binding residues and non-binding residues across the dataset for the ProtBert and SaProt models trained with equal positive and negative sampling. We find the models trained with equal positive and negative sampling have more interpretable attention maps despite a decreased performance on the LIT-PCBA benchmark compared to models trained with more negative samplings. We compare the attention of all the models in Figure S2, visualizing three otherwise identical models trained with different initial random seeds. Interestingly, the PLM with the worst performance on the LIT-PCBA benchmark, ProtBert, shows the most attention to binding site residues relative to its attention to non-binding site residues. The structure-aware PLM, SaProt, has two seeds that attend to binding residues more than non-binding residues across most of the attention heads and one seed that pays very little attention to the binding residues. The SaProt seed that has the least attention for binding residues as compared to non-binding residues performs the best on the LIT-PCBA benchmark. Across all PLMs, there is a large variance in the attention to binding residues as the models initial random seed is changed.
(a) ProtBert
(b) SaProt
(c) ProtBert-1to1
(d) SaProt-1to1
@@
Figure S2: Comparing the average attention weight of binding and non-binding residues on our set of 109 single-chain protein-ligand binding structures after training on the MERGED Dataset. The horizontal line indicates the average across the proteins. Each row is a different random seed and each column is a different PLM or different training regime, where ‘1to1’ indicates that a 1:1 positive to negative sampling ratio was used during training.
@@
We visualize the learned aggregation layers attention heads on the protein-ligand structures for both ProtBert and SaProt models in Figure 3 with additional visualizations provided in Appendix E and on our github.
Appendix E Structural visualizations

### v2

@@ 2.5 Pre-training to predict DTIs improves property prediction
To benchmark the usefulness of DTI co-embeddings for marginal property prediction, e.g., predicting the properties of a compound only from its molecular graph, we computed SPRINT DTI co-embeddings for several drug-like compounds [36] and natural products [37]. Concatenating the SPRINT-sm molecule co-embedding to a Morgan fingerprint consistently outperformed an equivalently sized neural network using only the Morgan fingerprint as input (Table 5). However, we observe that using only the SPRINT embedding in these tasks degrades performance, suggesting that SPRINT embeddings can synergistically enhance traditional fingerprints.
@@
3 Methods
In addition to classifying drug-target binding, SPRINT is capable of predicting binding affinity without changing learned portions of the model. By replacing the cosine similarity with a dot-product and removing the final sigmoid, we can leverage SPRINT as a binding affinity predictor (Equation 2):
$p\hat{K}=\left({Z_{d}}\cdot{Z_{t}}\right)$ (2)
where $p\hat{K}$ is the predicted $pK=-log_{10}(K_{d})$ to simplify training of the model on affinity values. Binding affinity regression models are trained to minimize mean square error loss against ground truth binding affinity by updating the weights of $C_{d}$ and $C_{t}$.
3.1 Data
We evaluate the performance of the SPRINT DTI prediction architecture on two tasks: interaction prediction and binding affinity prediction. SPRINT models trained for interaction prediction predict 1 if a drug and target pair interact and 0 otherwise. We utilize the same interaction prediction datasets as [2]: DAVIS, BIOSNAP, and BindingDB. DAVIS[40] and BindingDB[41] are both composed of drug-target paired with experimentally annotated dissociation constants, $K_{d}$. We utilize the same thresholds as [2] to set the binary labels: $K_{d}<30$ are labeled as interacting and $K_{d}\geq 30$ are non-interacting. DAVIS consists of only 2086 interactions, BindingDB has 12668, and BIOSNAP has 19238 DTIs. The BIOSNAP[42] data originally consisted of only interacting drug-target pairs, so negative pairs are created by randomly sampling an equal number of drug-target pairs. The Unseen Drugs and Unseen Targets datasets are different splits of BIOSNAP from [43] to elucidate the zero-shot performance of DTI methods. The Unseen Targets dataset is constructed by randomly selecting 20% of the proteins from the full BIOSNAP dataset and assigning all interactions involving these proteins to the test set, while the remaining interactions are used for training. The Unseen Drugs dataset is constructed in an identical manner, but using the drugs rather than the targets.
All dataset splits are the same as [43], 70/10/20 for training, validation, and test, respectively. The training split is artificially subsampled to have an equal number of positive and negative interactions. To evaluate the scaling of the model on more training tokens, we utilize the “MERGED” dataset[16]. This dataset is composed of data from PubChem [17] (98.31%), BindingDB [18] (1.17%), and ChEMBL [19] (0.52%). We use the activity label of the dataset for our interaction label providing 929,656 positive pairs and 83,703,190 negative pairs. We split the data into train, validation, and test splits such that targets in train and validation have at most 90% sequence homology, and targets in train and test have at most 70% sequence homology (splits are further detailed in Appendix C).
We use the Therapeutic Data Commons (TDC)[29] BindingDB_Patent dataset for binding affinity prediction. Each drug-target interaction is labeled with an experimentally determined $IC_{50}$. The BindingDB_Patent dataset splits train and test with a temporal split on the patent date, everything before 2019 is included in the train and validation split, while everything 2019 and after is in the test set. Temporal splits are still prone to data leakage as new drugs are commonly developed for established protein targets as well as drugs being reused for new targets [44]. The TDC hosts a leaderboard for this dataset to continuously evaluate new methods on the same training and test splits. We reserve a random 20% of the training set as a validation set.
Dataset unique drugs unique targets positive interactions negative interactions
DAVIS 68 379 1506 9597
BIOSNAP 4510 2181 13836 13647
BindingDB 10665 1413 9166 23435
MERGED 3529623 11958 929656 83703190
BindingDB_Patent 140746 477 N/A N/A
Table 4: Statistics of DTI datasets
4 Discussion
@@
Vector-based screens are extremely fast, enabling DTI prediction in regimes that would be impossible with structure-based approaches. We propose SPRINT, which improves on prior work using multi-head attention pooling that scales favorably as we increase the number of DTI training tokens. We show that SPRINT sets a new SOTA for DTI classification, virtual screening, and matches the performance of the top method on the TDC BindingDB_Patent leaderboard (Otter-Knowledge-Ensemble[30]) for binding affinity prediction, without the use of explicit knowledge graphs of protein-ligand interactions. We compare SPRINT using the ESM2 embeddings to their ensemble model using refined protein and ligand representations starting from the same representations as SPRINT. The SPRINT model achieves better binding affinity prediction than their four model ensemble which highlights the value of the attention pooling for the protein and the co-embedding space learned by the model. We demonstrate that SPRINT can perform virtual screening at pan-species proteome scales, e.g., for antimicrobial activities (Fig. 6). Lastly, we find that predicting DTIs via co-embedding is an effective pre-training strategy that enhances simple molecular property prediction (Table 5).
We also show that structure-aware PLMs like SaProt can confer performance gains in virtual screening. Interestingly, we find that the SPRINT models that perform best on the LIT-PCBA virtual screening benchmark, with increased negative sampling, have the least interpretable attention maps. We hypothesize that equal weighting of positive and negative drug-target pairs helps the model learn about residues that interact while increasing the amount of negatives dilutes the information of the positive examples. We compare the virtual screening performance of SPRINT on the second CACHE challenge to the second place computational method which utilizes DeepDocking. Notably, methods using DeepDocking have consistently placed in the top two for the three CACHE challenges in which final results are public, demonstrating its performance for finding hits from large libraries. SPRINT enables direct querying of the entire ENAMINE database, foregoing the inital random sample which can significantly bias the subsequent rounds of DeepDocking. After performing only 1/6th of the docking, SPRINT finds almost three times the number of high-scoring molecular scaffolds compared to gnina DeepDocking. SPRINT’s top molecules can be a useful starting point for DeepDocking to ensure that chemical space is more effectively explored.
@@
We envision the SPRINT framework and training task as a useful benchmarking tool for protein and molecule encoders. Future work will evaluate other structure-aware PLMs such as MULAN or S-PLM [45, 46], and other pre-trained molecule encoders like UniMol [47, 48], in the SPRINT framework for DTI prediction. Additional work to interpret the aggregated protein features through methods like InterPLM[49] could provide insight into the virtual screening predictions. We anticipate this technology will democratize virtual screening by enabling the discovery of lead molecules with a fraction of the compute cost required by comparable, structure-based, virtual screening methods.
[…]
Supplementary information
Additional information provided in the appendices.
[…]
Acknowledgements
The authors would like to thank Ian Dunn for providing the filtered dataset used for the CACHE screen in addition to the results of their virtual screening method for comparison.
Please refer to Journal-level guidance for any specific requirements.
Declarations
Funding
A.M. and D.K. are supported by the R35GM140753 from the National Institutes of Health. A.A. and H.M. are supported by the United States Department of Agriculture, Agriculture Research Service (USDA-ARS) and by grants 5R01GM107550-10 and 1U24DK133658-01. C.E. and E.X. are supported by the R01GM140467 from the National Institutes of Health, and the 2040381 from the National Science Foundation. M.D. was supported by the F31LM014194 from the National Institutes of Health.
Competing interests
The authors declare no competing interests.
Data availability
DAVIS, BIOSNAP, BindingDB, and MERGED DTI classification datasets are available through our github: https://github.com/abhinadduri/panspecies-dti. The BindingDB DTI DG dataset for the TDC Leaderboard is available through the Therapeutics Data Commons: https://tdcommons.ai/. The Lit-PCBA evaluation benchmark is available through the original authors https://drugdesign.unistra.fr/LIT-PCBA/. The filtered Enamine REAL dataset used for the second CACHE challenge is too large for upload, but can be obtained from the authors upon reasonable request.
Code availability
All code to reproduce our results is available on our github repository: https://github.com/abhinadduri/panspecies-dti.
Author contribution
A.M.: conceptualization, methodology, writing-original draft, software, investigation, validation A.A.: conceptualization, methodology, writing-original draft, software, investigation, validation C.E.: conceptualization, methodology, software, writing-original draft M.D.: software, visualization, writing-review&editing E.P.: supervision, funding acquisition H.M.: supervision, funding acquisition D.K.: conceptualization, resources, supervision, funding acquisition, writing-review&editing
@@ Appendix A SPRINT recovers known mechanisms of action
Our pan-species protein dataset is comprised of all predicted protein sequences in reference genomes from NCBI within the taxons bacteria, fungus, and human. Each taxon contained 3,379,854, 775,477, and 136,194 protein sequences respectively. We gathered a list of 3,112 natural products [37], as they are known to have a high prior likelihood for antimicrobial activity [50] and are out-of-distribution relative to our MERGED training dataset. We then co-visualized several antimicrobials and drugs with their known protein targets across microbial proteomes (Fig. 6), recovering several known mechanisms of action. The dataset is available to query at https://bit.ly/colab-screen.
Figure 6: UMAP visualization of the binding co-embedding space of drug-like small molecules and their protein targets across bacterial, fungal, and human proteomes. We see that antimicrobial compounds co-localize with regions of the shared latent space that contain human, bacterial, and fungal proteomes.
@@ Appendix B SPRINT co-embeddings improve property prediction
Current antimicrobial and toxicity screening approaches are often formulated as molecular property predictors, framing antimicrobial activity and toxicity to humans as inherent properties of drug molecules [51, 52, 53, 54, 36, 55]. Our results demonstrate that augmenting Morgan fingerprints with SPRINT-sm ligand embeddings consistently outperformed an equivalently sized neural network using Morgan fingerprints alone, when evaluated on both an antibacterial activity dataset [37] and a toxicity dataset [36] (Table 5). We hypothesize that vectorizing the DTI space allows property prediction methods to leverage information about target neighborhoods around a drug, enhancing performance and offering mechanistic explanations for these properties based on likely binding partners. The embeddings from the deeper SPRINT model consistently performed worse than those from the shallow SPRINT-sm model, suggesting that shallow transformations of the Morgan fingerprint work best in this setting. The standalone SPRINT embeddings achieved substantially lower performance than their concatenated counterparts, indicating that SPRINT embeddings may capture complementary molecular features to traditional fingerprints.
Table 5: F1 scores for MLP classification models applied to molecule embedding strategies (mean $\pm$ std). Models trained on Morgan fingerprints used a larger hidden size to match the size of models trained on fingerprints concatenated with embeddings.
@@
Morgan Fingerprint 0.740 $\pm$ 0.027 0.720 $\pm$ 0.008
SPRINT-sm Embedding 0.687 $\pm$ 0.012 0.656 $\pm$ 0.015
Morgan Fingerprint + SPRINT-sm Embedding 0.749 $\pm$ 0.016 0.735 $\pm$ 0.006
SPRINT Embedding 0.614 $\pm$ 0.027 0.631 $\pm$ 0.023
Morgan Fingerprint + SPRINT Embedding 0.722 $\pm$ 0.027 0.701 $\pm$ 0.018
@@ Appendix C Training details
Specifically, the loss $\mathcal{L}$ is written as
@@
where the protein, $T^{i}$, and drug, $D^{i}$, have been mapped to the SPRINT co-embedding space as $Z_{t}^{i}$ and $Z_{d}^{i}$, respectively. $Y_{i}\in\{0,1\}$ is a ground-truth label with value 1 if $T^{i}$ and $D^{i}$ are binders or 0 if they are non-binders. The pre-sigmoid scalar value, $\alpha$, is used to expand the range of cosine-similarity to the domain of the sigmoid. We set $\alpha$ to 5. Models are trained for binary classification using standard, supervised cross-entropy loss.
Binding affinity regression models use the same architecture to learn the co-embedding space, but the final, non-learned components are removed. The sigmoid and pre-sigmoid scalar value are removed and the cosine similarity is replaced with a dot-product (Equation 2). This allows the magnitude of drug and target embeddings to impact the prediction as well as increasing the range to all real numbers. Regression models are trained with mean squared error loss.
ConPLex models are trained with the hyperparameters used in the original paper [2]. Our Attention Pooling models are trained with a learning rate of $1\times 10^{-5}$ and a dropout value of 0.05 for 250 epochs, keeping all other hyperparameters the same as ConPLex training. We set weight decay to 0.01 for the binding affinity regression model following hyperparameter tuning on the validation set. The model checkpoint with the highest validation AUPR for classification models and MSE for regression models during training is evaluated on the test set (Table 1 and 3).
To enable efficient training on the MERGED dataset, we featurize the unique proteins and molecules before training, storing their representations in memory-mapped files for quick retrieval using the Lightning Memory-Mapped Database Manager (LMDB) library. To address data imbalance in the binding data, for each epoch, we train using all of the drug-target binding pairs, and subsample an equivalent number of non-binding pairs without replacement. We observed that models trained with more negatives than positives (at a 3:1 ratio), achieved better virtual screening performance, but had less interpretable attention patterns (Tables 1, 6). Models are trained for 20 epochs. All other hyperparameters were kept the same.
Table 6: LIT-PCBA evaluation ablation of negative sampling. ‘1:1’ indicates equal sampling of positive and negative examples during training and ‘3:1’ indicates the preferred model training with 3 negatives sampled for every positive example.
@@
Appendix D High scoring molecules screened with SPRINT for NSP13 Helicase
We show the 2D depiction of the highest scoring molecules, according to Gnina’s CNN VS, from SPRINT’s screen of Enamine REAL in Figure 7. We see that the molecules represent diverse scaffolds.
Figure 7: 2D depiction of the high-scoring (CNN VS $>$ 6) molecules found by SPRINT for NSP13 Helicase.
Appendix E Investigating the learned aggregation layer
@@
We first analyzed the attention patterns of each head in the learned aggregation layer to determine if any of the heads were selective for binding residues. We calculate the attention scores for each residue in the protein and compute the mean attention value for the binding residues and the non-binding residues. Figure 2 shows the average weight of the binding residues and non-binding residues across the dataset for the ProtBert and SaProt models trained with equal positive and negative sampling. We find the models trained with equal positive and negative sampling have more interpretable attention maps despite a decreased performance on the LIT-PCBA benchmark compared to models trained with more negative samplings. We compare the attention of all the models in Figure 8, visualizing three otherwise identical models trained with different initial random seeds. Interestingly, the PLM with the worst performance on the LIT-PCBA benchmark, ProtBert, shows the most attention to binding site residues relative to its attention to non-binding site residues. The structure-aware PLM, SaProt, has two seeds that attend to binding residues more than non-binding residues across most of the attention heads and one seed that pays very little attention to the binding residues. The SaProt seed that has the least attention for binding residues as compared to non-binding residues performs the best on the LIT-PCBA benchmark. Across all PLMs, there is a large variance in the attention to binding residues as the models initial random seed is changed.
@@
Figure 8: Comparing the average attention weight of binding and non-binding residues on our set of 109 single-chain protein-ligand binding structures after training on the MERGED Dataset. The horizontal line indicates the average across the proteins. Each row is a different random seed and each column is a different PLM or different training regime, where ‘1to1’ indicates that a 1:1 positive to negative sampling ratio was used during training.
@@
We visualize the learned aggregation layers attention heads on the protein-ligand structures for both ProtBert and SaProt models in Figure 3 with additional visualizations provided in Appendix F and on our github.
Appendix F Structural visualizations